OpenStudy (anonymous):

What is the 9th term of the geometric sequence 4, –20, 100, …?

OpenStudy (anonymous):

Multiplying by -5. 4 x -5 = -20 -20 x -5 = 100 etc. the third term is 100, so keep going :) (hint: answer will be positive)

OpenStudy (anonymous):

so u just told me what i already know?